if a=xy^(p-1),b=xy^(q-1) and c=xy^(r-1).prove that a^(o-r).b^(p).q^(p-q)=1

Grade:9

1 Answers

Riddhish Bhalodia
8 years ago
I am pretty sure the you made a slight mistake in typing the question
you would have to prove
a^{q-r}b^{r-p}c^{p-q} = 1
so to do that just substitute the values of a b and c
we get
LHS = x^{q-r+r-p+p-q}y^{(p-1)(q-r) + (q-1)(r-p) + (r-1)(p-q)}
LHS = x^{0}y^{pq-pr-q+r+qr-pq-r+p+pr-qr-p+q} = x^0y^0 = 1 = RHS
hence proved